The dimensions of a rectangular garden were 3 m x 10 m. When both dimensions were increased by the same amount, the area of the garden doubled. Find the dimensions of the new garden. plz help
Answer by mananth(16946)

The dimensions of a rectangular garden were 3 m x 10 m. When both dimensions were increased by the same amount, the area of the garden doubled. Find the dimensions of the new garden.
The area of the garden is 3*10 = 30 m^2
Let the dimensions be increased by x m
new dimensions
(3+x) & (10+x)
both dimensions were increased by the same amount, the area of the garden doubled.
(3+x)(10+x) = 60
x^2+13x+30=60
x^2+13x-30=0
x^2+15x-2x-30=0
x(x+15)-2(x+15)=0
(x+15)(x-2)=0
x=-15 OR x=2
take positive root
x=2
the new dimensions are 5m & 12 m